# FoundationGrasp: Generalizable Task-Oriented Grasping with Foundation Models

### v1

@@ IV-B Data Generation
Geometric descriptions focus on describing two aspects: (1) the geometric structure of $c_{i}$ and (2) the correlations between geometric structures (form) and $t_{j}$ (function). To address the former, we design an object-to-part (O2P) prompt set, with each prompt querying the geometric structure of $c_{i}$ (e.g., “Describe what parts the household object $c_{i}$ is composed of:”). In parallel, for each task $t_{j}$, we design a task-to-part (T2P) prompt set, with each prompt querying geometric structures that support $t_{j}$ (e.g., “Describe what geometric primitives on a household object can be used to $t_{j}$:”). We regularize the outputs of the LLM to bullets of parts or primitive shapes. This regularization strategy reduces the difficulty of grounding geometric descriptions to 3D representations during task-oriented grasp evaluation.
@@
Web-Based Image Retrieval  Previous works in 3D scene representation have demonstrated that visual-language features distilled from 2D foundation models facilitate geometric understanding of 3D world [36, 37]. Following the generation of geometric descriptions, we utilize the Internet as a large-scale database to retrieve visual data. Given that each instance in the TaskGrasp dataset is captured from only 2-3 viewpoints, we augment the dataset by crowdsourcing images from the Internet. Specifically, for each object instance $o_{k}$ in the dataset, we first query Google Cloud Vision API with its original images in the dataset, retrieving candidate images with respective confidence scores. Each candidate image contains an object instance sharing similar visual characteristics and geometric structures. Subsequently, we employ the Segment Anything Model (SAM) [38] for background removal. To guarantee the view diversity, we utilize CLIP [39] visual encoder to extract the embedding vector of each candidate image and compute the cosine similarities between them. Those with high similarities are excluded. Finally, a set of $N_{I}$ images (original + retrieved images) for $o_{k}$, denoted as $I_{o_{k}}$, is selected based on confidence scores. Examples of retrieved images are available in the appendix. Note that the goal here is not to perform accurate 3D reconstruction of $o_{k}$. Thus, the instances in the retrieved images do not need to match $o_{k}$ precisely. This image retrieval strategy aims to provide visual information for constructing geometric priors of objects and tasks in the later stage. During inference, the robot utilizes $I_{o}$ captured by the onboard camera.
@@ IV-C Multi-Modal Feature Representation
Object and Grasp Encoder  To model the relative spatial relationship between the gripper and the object, the first step involves transforming $g\in SE(3)$ and $X_{o}\in\mathbb{R}^{N\times 3}$ into the same space. This is achieved by approximating $g$ with six control points[19], represented as $X_{g}\in\mathbb{R}^{6\times 3}$ in the object frame (shown in Figure 2). Then, $X_{o}$ and $X_{g}$ are jointly embedded. Specifically, $X_{o}$ and $X_{g}$ are first concatenated to form $X_{o+g}\in\mathbb{R}^{(N+6)\times 3}$. A binary feature vector is appended to indicate whether a point belongs to $X_{o}$ or $X_{g}$. Subsequently, a PointNet++[40] like encoder (denoted as PN++) extracts localized features from $X_{o+g}$, which associates features to a reduced number of $N^{-}$ points ($N^{-}<N$):
@@
where point embeddings $F_{X}\in\mathbb{R}^{N^{-}\times 1024}$. We choose to maintain a degree of locality in the latent space, as opposed to relying on a single global embedding, to preserve the localized geometric information. Such information is necessary for understanding the relationship between geometric structures, objects, and tasks, as discussed in Section IV-B. The point encoder consists of three set abstraction (SA) layers with multi-scale grouping (MSG). A Multi-Layer Perceptron (MLP) is then applied for dimension reduction to output $\tilde{F}_{X}\in\mathbb{R}^{N^{-}\times 256}$.
@@ IV-D Task-Oriented Grasp Evaluation
Semantic Branch  $\textnormal{TGE}_{\textnormal{sem}}$ aims to linguistically connect novel objects and tasks with familiar ones using semantic descriptions. The problem is analogous to machine translation in natural language processing (NLP), which aims to transform language sequences from the source to the target domain. In our context, the robot initially does not understand novel elements in $L$. $\textnormal{TGE}_{\textnormal{sem}}$ translates them to familiar ones using semantic descriptions. Inspired by this similarity, we design $\textnormal{TGE}_{\textnormal{sem}}$ based on the Transformer decoder, an effective architecture for handling sequence-to-sequence problems.
@@
where $\tilde{F}_{I}^{geo}\in\mathbb{R}^{hw\times 256}$ ($\tilde{F}_{I,c}^{geo}$ or $\tilde{F}_{I,t}^{geo}$ ). The architecture of $\textnormal{TGE}_{\textnormal{geo}}$ is depicted in Figure 3 (left). Similar to the semantic branch, the geometric branch is divided into two object and task knowledge incorporation, which correspond to $\phi_{c}^{geo}$ and $\phi_{t}^{geo}$ in $\textnormal{TGE}_{\textnormal{geo}}$, respectively. Each layer incorporates geometric features $\tilde{F}^{geo}_{I}$ into $\tilde{F}_{X}$:
@@
where MHA consists of $h_{geo}$ heads. The computational procedure of cross-attention fusion is similar to Equation 6. The self-attention rectification operates similarly to cross-attention fusion, with the distinction that query, key, and value tensors are all projections of $\tilde{F}_{X}^{geo}$. The design choice, where a rectification layer follows a fusion layer, is based on the consideration that a self-attention layer performs a permutation-invariant transformation to rectify geometric features that are not properly aligned[42]. Finally, $\tilde{F}_{X}^{geo}$ is processed by an SA layer with global pooling to output the point embedding $\overline{F}_{X}^{geo}\in\mathbb{R}^{300}$ contextualized with geometric knowledge.
@@
Loss Function  The binary cross-entroy loss between $S$ and the ground truth label $S_{gt}$ is computed as below:
@@ V-A Perception Experiments
Baselines  We compare FoundationGrasp to the following baselines that are representative of existing grasp synthesis methods:
@@
  • SGN is an adapted version of Semantic Grasp Network proposed in [30]. It learns object-task-grasp relationships solely from training data without incorporating external knowledge.
@@
  • GCNGrasp [5] is the state-of-the-art graph-based TOG method. It incorporates a semantic KG connecting a pre-defined set of objects and tasks. In our implementation, given a novel object class or task out of the graph, we connect the node to its nearest neighbor in the graph by measuring the cosine similarity of their ConceptNet pre-trained word embeddings.
@@
Figure 4: Test objects in real-robot experiments cover commonly used accessories, kitchen utensils, and mechanic tools.
@@
Method Held-out Instance Performance Held-out Class Performance Held-out Task Performance
Instance mAP Class mAP Task mAP Instance mAP Class mAP Task mAP Instance mAP Class mAP Task mAP
TAG 59.75 60.28 54.76 59.32 58.73 52.72 59.06 58.24 52.37
SGN 78.51 75.08 68.80 74.20 72.95 62.55 75.17 71.59 63.35
SGN+we 79.74 77.91 74.36 77.21 75.51 63.73 78.06 74.49 70.55
GCNGrasp 83.24 82.02 73.71 72.92 72.45 67.58 57.48 47.21 33.17
GraspGPT 80.25 77.94 76.10 79.70 77.88 72.84 79.32 76.90 72.34
FoundationGrasp 93.17 91.55 89.39 88.74 86.09 84.33 89.01 87.36 82.70
TABLE III: Quantitative results of perception experiments
@@ V-B Real-Robot Experiments
Task-Oriented Grasping Experiments  The experiment platform comprises a 7 DoF Kinova Gen3 robotic arm equipped with a Robotiq 2F-85 parallel jaw gripper and an Intel RealSense D435 RGB-D camera with eye-in-hand calibration. Shown in Figure 4, we collect test objects from the YCB dataset and our laboratory, covering commonly used accessories, kitchen utensils, and mechanic tools. During inference, SAM [38] is first applied to extract the test object from the point cloud captured from a single view. Contact-GraspNet [21] then samples $N_{g}^{+}$ grasp candidates for evaluation. Compared to the multi-view setting in perception experiments, the single-view setup is used in real-robot experiments since a robot, in most cases, only has a partial observation of the object.
Task-Oriented Manipulation Experiments  To support the hypothesis that FoundationGrasp facilitates downstream tool manipulation, we combine FoundationGrasp with pre-programmed manipulation skills and conduct task-oriented manipulation experiments. Presented in Table V, we conduct manipulation experiments on nine household tasks selected from the household chores list in [44]. Since manipulation policy learning is not the focus of this paper, we design rule-based heuristics to determine the operating direction of the tool object and compute the effect point on the receptacle object. Manipulation trajectories are generated by merging sequential motion primitives.
Metrics  The task-oriented grasping pipeline is divided into three stages: perception, planning, and action. We report the statistics of each stage separately, providing details for failure analysis. A trial is considered successful if the test object is grasped by the robot in a task-oriented manner given a language instruction. For task-oriented manipulation experiments, we record the success of task-oriented grasping and manipulation, respectively.
@@ V-C Implementation Details
Training and inference are conducted on a desktop PC with a single NVIDIA RTX 3090 GPU. All the baselines are re-trained on the LaViA-TaskGrasp dataset for fair comparison. FoundationGrasp is implemented using Pytorch 1.11.0 with CUDA 11.3. We use an Adam [45] optimizer with a weight decay of 0.0001. The learning rate is set to 0.0001 initially and adjusted with a LambdaLR scheduler as in [5, 14]. To increase the robustness to unseen object poses, input point clouds are augmented with random scaling, rotation, jitter, and dropout. Each point cloud is then downsampled to 4096 points and fed into FoundationGrasp. The model is trained for 50 epochs with a batch size of 32.
For the choice of foundation models, we use OpenAI GPT-4 to generate language descriptions and OpenAI CLIP implemented by OpenCLIP as the VLM. For the large-pretrained language encoder, we choose the Google pre-trained BERT-Base model provided by Hugging Face. All the foundation models are frozen during training. More details on model training and design can be found in the appendix.
Figure 5: Qualitative results of perception experiments. GT denotes ground truth task-oriented grasp poses. FoundationGrasp, GCNGrasp, and GraspGPT are evaluated under three held-out settings. Top-5 predictions of each method are visualized. Grasp poses are colored by their confidence scores (green is higher).
@@ VI-A Results of Perception Experiments
The quantitative results of perception experiments are reported in Table III. For three held-out settings, TAG establishes a lower performance bound for all methods, achieving approximate mAPs of 50-60%. The seemingly high mAPs by random guessing is because the distribution of positive and negative samples in the LaViA-TaskGrasp dataset is about even. In the case where the dataset is annotated with more negative samples, mAPs are expected to be much lower. SGN surpasses TAG’s performance by over 10% since it considers the task constraints on object grasping. Moreover, ConceptNet word embeddings capture the semantic similarity between relevant concepts in the embedding space. Consequently, SGN+we achieves a marginal performance boost compared to SGN.
GCNGrasp performs decently well, with mAP scores of 83.24%, 82.02%, and 73.71%, when dealing with intra-class variances. It can be attributed to the fact that the semantic KG pre-defines relations between all classes and tasks in the dataset. Nonetheless, GCNGrasp experiences significant performance drops when generalizing to novel classes and tasks out of KG due to its inherent closed-world nature. Such a limitation becomes especially evident in the held-out task setting, where GCNGrasp even falls behind TAG by 1.58%, 11.03%, and 19.20%, and SGN by 17.69%, 24.38%, and 30.18% across three evaluation metrics. It also suggests that ConceptNet word embeddings are more suitable for capturing linguistic relations between objects than tasks.
@@
FoundationGrasp and GraspGPT outperform TAG, SGN, and GCNGrasp in both the held-out class and task settings, demonstrating the superiority of the open-ended design. Due to the incorporation of geometric knowledge, FoundationGrasp further improves the performance of GraspGPT by approximately 10%. Overall, FoundationGrasp performs best in all held-out settings. Regarding running time, it takes less than two seconds for FoundationGrasp to evaluate 25 grasp candidates on a single RTX 3090. A qualitative comparison between all methods is illustrated in Figure 5. Notably, even though baselines may assign high confidence scores to their predictions, they often fail to accurately localize the task-oriented regions. FoundationGrasp also exhibits commonsense reasoning capability to some degree. As is depicted in Figure 6, FoundationGrasp rejects all grasp candidates (i.e., assign low confidence scores) under conditions where the object is incompatible with the intended task, or there is a discrepancy between the target and actual object classes.
VI-B Results of Real-Robot Experiments
Task-Oriented Grasping Experiments  The result of task-oriented grasping experiments is summarized in Table IV. 100 trials are conducted for each held-out setting, with 10 per instance, class, or task. Despite capturing object point clouds from a single viewpoint, FoundationGrasp consistently achieves a perception success rate exceeding 80% across all settings. Similar to earlier observations in perception experiments, GCNGrasp suffers when generalizing to novel classes and tasks, while GraspGPT and FoundationGrasp perform consistently better. The improvement observed with FoundationGrasp over GraspGPT indicates the necessity of incorporating both semantic and geometric knowledge. FoundationGrasp experiences a significant performance drop from the perception to the action stage. Three typical failure modes are (1) unstable grasp candidates generated by the task-agnostic grasp sampler, (2) motion planning failures, and (3) false negative evaluations by FoundationGrasp (e.g., rejecting all valid candidates). Concerning the third failure mode, the training stage of FoundationGrasp can be conceptualized as constructing an implicit knowledge base in the latent space with a set of objects and tasks. During the inference stage, FoundationGrasp aims to build connections between existing elements in the knowledge base and novel ones outside. However, if novel elements are substantially divergent from the knowledge base, FoundationGrasp tends to make false negative predictions. As illustrated in Figure 8, in the case of instructing the robot to screw with a power drill, object class power drill shares minimal semantic or geometric similarity to existing elements in the knowledge base, leading to the rejection of all valid candidates. Nevertheless, FoundationGrasp achieves the highest grasping success rate, averaging 72.67% over three held-out settings.
@@ VI-B Results of Real-Robot Experiments
TABLE IV: Quantitative results of task-oriented grasping experiments
@@
Figure 8: Since the novel object class “Power Drill” shares minimal similarity to familiar elements in the implicit knowledge base, FoundationGrasp makes false positive predictions by rejecting all valid candidates. The colors of novel elements denote the level of similarity to the “Mug”.

### v2

@@ IV-B Knowledge Generation
Geometric descriptions focus on two aspects: (1) the geometric structure of $c_{i}$ and (2) the correlations between geometric structures (form) and $t_{j}$ (function). To address the former, we design an object-to-part (O2P) prompt set, with each prompt querying the geometric structure of $c_{i}$ (e.g., “Describe what parts the household object $c_{i}$ is composed of:”). In parallel, for each task $t_{j}$, we design a task-to-part (T2P) prompt set, with each prompt querying geometric structures that support $t_{j}$ (e.g., “Describe what geometric primitives on a household object can be used to $t_{j}$:”). We regularize the outputs of the LLM to bullets of parts/primitive shapes. This regularization strategy reduces the difficulty of grounding geometric descriptions to 3D representations during task-oriented grasp evaluation and meanwhile prevents the LLM from hallucinating.
@@
Web-Based Image Retrieval  Previous works in 3D scene representation have demonstrated that visual-language features distilled from 2D foundation models facilitate geometric understanding of 3D world [41, 42]. Following the generation of language descriptions, we augment the TaskGrasp dataset by utilizing the Internet as a large-scale database to retrieve visual data. Each instance in the original TaskGrasp dataset is captured from only 2-3 viewpoints. Therefore, for each object instance $o_{k}$ in the dataset, we first query Google Cloud Vision API with its original images in the dataset, retrieving candidate images with respective confidence scores. Each candidate image contains an object instance sharing similar visual characteristics and geometric structures. Subsequently, we employ the Segment Anything Model (SAM) [43] for background removal. To guarantee the view diversity, we utilize CLIP [44] visual encoder to extract the embedding vector of each candidate image and compute the cosine similarities between them. Those with high similarities are excluded. Finally, a set of $N_{I}$ images (original + retrieved images) for $o_{k}$, denoted as $I_{o_{k}}$, is selected based on confidence scores. Examples of retrieved images are available in the appendix. Note that the goal here is not to perform accurate 3D reconstruction of $o_{k}$. Thus, the instances in the retrieved images do not need to match $o_{k}$ precisely. This image retrieval strategy aims to provide visual information for constructing geometric priors of objects and tasks in the later stage. During inference, the robot utilizes $I_{o}$ captured by the onboard camera.
@@ IV-C Multi-Modal Feature Representation
Object and Grasp Encoder  To model the relative spatial relationship between the gripper and the object, the first step involves transforming $g\in SE(3)$ and $X_{o}\in\mathbb{R}^{N\times 3}$ into the same space. Following the variational grasp generator 6-DoF GraspNet [24], the first step is achieved by approximating $g$ with six control points, represented as $X_{g}\in\mathbb{R}^{6\times 3}$ in the object frame (shown in Figure 2). Then, $X_{o}$ and $X_{g}$ are jointly embedded. Specifically, $X_{o}$ and $X_{g}$ are first concatenated to form $X_{o+g}\in\mathbb{R}^{(N+6)\times 3}$. A binary feature vector is appended to indicate whether a point belongs to $X_{o}$ or $X_{g}$. Subsequently, a point encoder (denoted as PN++) extracts localized features from $X_{o+g}$, which associates features to a reduced number of $N^{-}$ points ($N^{-}<N$):
@@
where point embeddings $F_{X}\in\mathbb{R}^{N^{-}\times 1024}$. PN++ is adopted from PointNet++[45], a deep learning architecture designed for directly processing 3D point cloud data without losing the inherent structure and details. We choose to maintain a degree of locality in the latent space, as opposed to relying on a single global embedding, to preserve the localized geometric information. Such information is necessary for understanding the relationship between geometric structures, objects, and tasks, as discussed in Section IV-B. The point encoder consists of three set abstraction (SA) layers with multi-scale grouping (MSG). A Multi-Layer Perceptron (MLP) is then applied for dimension reduction to output $\tilde{F}_{X}\in\mathbb{R}^{N^{-}\times 256}$.
@@ IV-D Task-Oriented Grasp Evaluation
Semantic Branch  $\textnormal{TGE}_{\textnormal{sem}}$ aims to linguistically connect novel objects and tasks with familiar ones using semantic descriptions. The problem is analogous to machine translation in natural language processing (NLP), which aims to transform language sequences from the source to the target domain. In our context, the robot initially cannot understand novel elements in $L$. $\textnormal{TGE}_{\textnormal{sem}}$ translates them to familiar ones using semantic descriptions. Inspired by this similarity, we design $\textnormal{TGE}_{\textnormal{sem}}$ based on the Transformer decoder, an effective architecture for handling sequence-to-sequence problems.
@@
where $\tilde{F}_{I}^{geo}\in\mathbb{R}^{hw\times 256}$ ($\tilde{F}_{I,c}^{geo}$ or $\tilde{F}_{I,t}^{geo}$ ). The architecture of $\textnormal{TGE}_{\textnormal{geo}}$ is depicted in Figure 3 (left). Similar to the semantic branch, the geometric branch is divided into object and task knowledge incorporation, which correspond to $\phi_{c}^{geo}$ and $\phi_{t}^{geo}$ in $\textnormal{TGE}_{\textnormal{geo}}$, respectively. Each layer incorporates geometric features $\tilde{F}^{geo}_{I}$ into $\tilde{F}_{X}$:
@@
where MHA consists of $h_{geo}$ heads. The computational procedure of cross-attention fusion is similar to Equation 6. The self-attention rectification operates similarly to cross-attention fusion, with the distinction that query, key, and value tensors are all projections of $\tilde{F}_{X}^{geo}$. The design choice, where a rectification layer follows a fusion layer, is based on the consideration that a self-attention layer performs a permutation-invariant transformation to rectify geometric knowledge features that are not properly aligned[47]. Finally, $\tilde{F}_{X}^{geo}$ is processed by an SA layer with global pooling to output the point embedding $\overline{F}_{X}^{geo}\in\mathbb{R}^{300}$ contextualized with geometric knowledge.
@@
Loss Function  The binary cross-entropy loss between $S$ and the ground truth label $S_{gt}$ is computed as below:
@@ V-A Perception Experiments
Baselines  We compare FoundationGrasp to the following baselines that are representative of existing methods:
@@
  • SGN is an adapted version of Semantic Grasp Network [34]. It learns object-task-grasp relationships solely from training data without incorporating external knowledge.
@@
  • GCNGrasp [10] is the state-of-the-art graph-based TOG method. It utilizes a semantic KG connecting a pre-defined set of objects and tasks. In our implementation, if an object class or task is not present in the graph, we connect the node to its nearest neighbor in the graph by measuring the cosine similarity of their ConceptNet pre-trained word embeddings. Specifically, we compare FoundationGrasp with two versions of GCNGrasp: the open-world version (GCNGrasp-ow) and the closed-world version (GCNGrasp-cw). The open-world version adheres to the same train/test split as other baselines, while the closed-world version assumes access to all (i.e., training and test) object classes and tasks in its KG, along with their ground truth relations. Although this assumption is impractical for real-world robotic applications, it allows us to highlight the performance gap between generalizing to elements within versus outside the pre-defined sets. By default, GCNGrasp refers to GCNGrasp-ow unless specified otherwise.
@@
Figure 4: Novel objects tested in real-robot experiments cover commonly used accessories, kitchen utensils, and mechanic tools.
@@ V-B Real-Robot Experiments
Single-Object Task-Oriented Grasping Experiments  The experiment platform consists of a 7-DoF Kinova Gen3 robotic arm equipped with a Robotiq 2F-85 parallel jaw gripper and an Intel RealSense D435 RGB-D camera with eye-in-hand calibration. As shown in Figure 4, we collect 42 test objects unseen during training from the YCB dataset and our laboratory, with variations in functionality, use case, geometry, and material. The test objects cover commonly used accessories, kitchen utensils, and mechanic tools. During each trial, a test object is randomly selected and placed on a flat-surface kitchen table with varying poses. The robot then receives a language instruction from the user specifying the target object and the desired task. During inference, SAM is first applied to extract the test object from the scene point cloud captured from a single view. Contact-GraspNet [26] then samples $N_{g}^{+}$ grasp candidates for evaluation. Finally, the robot picks the grasp candidate with the highest score for execution. Compared to the multi-view setting in perception experiments, the single-view setup is used in real-robot experiments since a robot, in most cases, only has a partial observation of the object.
Multi-Object Task-Oriented Grasping Experiments  The multi-object TOG experiments are conducted at three levels of clutter: (1) light clutter, where objects are spaced apart and clearly visible without occlusion in the camera view; (2) medium clutter, where light occlusions and contacts exist between objects with some functional parts not fully visible; (3) heavy clutter, where significant physical contacts and heavy occlusions occur between objects or functional parts. For each trial, a random set of 3 to 10 objects is selected, including distractor objects out of the test object set (e.g., apple, plate). Detection and segmentation are performed using GroundingDINO [49] and SAM. The rest of the setup remains consistent with that used in the single-object TOG experiments.
Task-Oriented Manipulation Experiments  To support the hypothesis that FoundationGrasp facilitates downstream tool manipulation, we combine FoundationGrasp with pre-programmed manipulation skills and conduct task-oriented manipulation experiments. Presented in Table VI, we conduct manipulation experiments on 9 household tasks selected from the household chores list [50]. Since manipulation policy learning is not the focus of this paper, we design rule-based heuristics to determine the operating direction of the tool object and compute the effect point on the receptacle object. Manipulation trajectories are generated by merging sequential motion primitives.
Metrics  The single-object TOG pipeline is divided into three stages: Perception, Planning, and Action. We report the statistics for each stage separately, providing details for failure analysis. 100 trials are conducted on the held-out instance, class, and task settings, with 10 per instance, class, or task. A trial is considered successful if the robot correctly detects TOG poses, plans the end effector to the top-ranked TOG pose, and physically executes the pose on the target object. Since there are no ground truth TOG labels for test objects, task compatibility is determined manually by the user. Stability is verified if the object is grasped and lifted for three seconds by the robot. In the multi-object TOG experiments, we additionally report the success of object grounding, referred to as “Grounding,” to indicate whether the robot correctly selects the target object. If multiple objects meet the language instruction criteria (e.g., two mugs for drinking), successfully executing TOG poses on any of them is considered a success. 15 trials are conducted at each clutter level, resulting in 45 trials in total. For task-oriented manipulation experiments, we record the success rates for both task-oriented grasping and manipulation. Each of the 9 household tasks is tested with 10 trials.
@@ V-C Implementation Details
Training and inference are conducted on a desktop PC with a single NVIDIA RTX 3090 GPU. All the baselines are re-trained and tested on the LaViA-TaskGrasp dataset with the same loss function for fair comparison. FoundationGrasp is implemented using Pytorch 1.11.0 with CUDA 11.3. We use an Adam [51] optimizer with a weight decay of 0.0001. The learning rate is set to 0.0001 initially and adjusted with a LambdaLR scheduler as in [10, 19]. To increase the robustness to unseen object poses, input point clouds are augmented with random scaling, rotation, jitter, and dropout. Each point cloud is then downsampled to 4096 points and fed into FoundationGrasp. The model is trained for 50 epochs with a batch size of 32.
For the choice of foundation models, we use OpenAI GPT-4 to generate language descriptions and OpenAI CLIP implemented by OpenCLIP as the VLM. For the large pre-trained language encoder, we choose the Google pre-trained BERT-Base model provided by Hugging Face. All the foundation models are frozen during training. To facilitate the training efficiency, the knowledge described in Section IV-B is pre-generated and pre-processed with selected foundation models before training. More details on model training and design can be found in the appendix.
For real-robot experiments, after determining the target grasp pose, we use MoveIt![52] for trajectory generation. Specifically, KDL[53] and RRT-Connect[54] are used for inverse kinematics computation and trajectory planning, respectively. Collision checking is performed using the FCL[55] package. Upon reaching the target grasp pose, pre-defined gripper aperture settings and maximum torque are executed by the controller of the Robotiq 2F-85 gripper.
Figure 5: Quantitative results of perception experiments
Figure 6: Qualitative results of perception experiments. GT denotes ground truth task-oriented grasp poses. Top-5 predictions of each method are visualized. Grasp poses are colored by their confidence scores (green is higher).
@@ VI-A Results of Perception Experiments
The quantitative results of perception experiments are visualized in Figure 5. For three held-out settings, TAG establishes a lower performance bound for all methods, achieving approximate mAPs of 50-60%. The seemingly high mAPs by random guessing is because the distribution of positive and negative samples in the LaViA-TaskGrasp dataset is about even. In the case where the dataset is annotated with more negative samples, mAPs are expected to be much lower. SGN surpasses TAG’s performance by over 10% since it considers the task constraints on object grasping. Moreover, ConceptNet word embeddings capture the semantic similarity between relevant concepts in the embedding space. Consequently, SGN+we achieves a marginal performance boost compared to SGN.
In the held-out instance setting, where no object classes and tasks are held out from the pre-defined KG, GCNGrasp-ow and GCNGrasp-cw are essentially the same, achieving mAP scores of 83.24%, 82.02%, and 73.71%. This is because the semantic KG pre-defines all the object classes, tasks, and their semantic relations within the dataset. However, the performance gap between GCNGrasp-ow and GCNGrasp-cw becomes evident in the other held-out settings. GCNGrasp-ow suffers when generalizing to novel classes and tasks outside the KG. It even falls behind TAG by 1.58%, 11.03%, and 19.20%, and behind SGN by 17.69%, 24.38%, and 30.18% across the three evaluation metrics. In contrast, since GCNGrasp-cw assumes access to all the object classes, tasks, and their ground truth relations in the KG, it performs consistently better. Comparing the two versions of GCNGrasp, the performance degrades by 6.43%, 4.43%, and 5.39% in the held-out class setting, and by 22.95%, 28.85%, and 42.94% in the held-out task setting. The significant degradation highlights the limitation of previous methods. The differences between the two held-out settings suggest that while the pre-trained word embeddings from ConceptNet can effectively capture linguistic relations between object classes, they are less effective at relating task concepts.
@@
With the open-ended knowledge from foundation models, FoundationGrasp and GraspGPT outperform TAG, SGN, and GCNGrasp in both held-out class and task settings. Even though FoundationGrasp does not assume access to all the objects and tasks in the dataset during training as GCNGrasp-cw does, it still outperforms GCNGrasp-cw. Due to the incorporation of geometric knowledge, FoundationGrasp further improves the performance of GraspGPT by approximately 10%. Overall, FoundationGrasp performs the best in all held-out settings. Regarding running time, it takes an average of 1.88s for FoundationGrasp to evaluate 25 grasp candidates per instance in the LaViA-TaskGrasp on a single RTX 3090. A qualitative comparison between all methods is illustrated in Figure 6. Note that even though baselines may assign high confidence scores to their predictions, they often fail to accurately localize the task-oriented regions. FoundationGrasp also exhibits commonsense reasoning capability to some degree. As depicted in Figure 7, FoundationGrasp rejects all grasp candidates (i.e., assigns low confidence scores) under conditions where the object is incompatible with the intended task, or there is a discrepancy between the target and actual object classes.
@@
TABLE III: Quantitative results of single-object task-oriented grasping experiments
Method Light Clutter Success Medium Clutter Success Heavy Clutter Success
Grounding Perception Planning Action Grounding Perception Planning Action Grounding Perception Planning Action
FoundationGrasp 14/15 13/15 12/15 12/15 80.00% 13/15 11/15 11/15 11/15 73.33% 10/15 9/15 8/15 8/15 53.33%
TABLE IV: Quantitative results of multi-object task-oriented grasping experiments
@@
Figure 8: Since the novel object class “Power Drill” shares minimal similarity to familiar elements in the implicit knowledge base, FoundationGrasp makes false predictions by rejecting all valid candidates. The colors of novel elements denote the level of similarity to the “Mug”.
Figure 9: Qualitative results of multi-object task-oriented grasping experiments
VI-B Results of Real-Robot Experiments
Single-Object Task-Oriented Grasping Experiments  The results of single-object task-oriented grasping experiments are summarized in Table III. Despite capturing object point clouds from a single viewpoint, FoundationGrasp consistently achieves a perception success rate exceeding 80% across all settings. Similar to earlier observations in perception experiments, GCNGrasp suffers when generalizing to novel classes and tasks, while GraspGPT and FoundationGrasp perform consistently better. The improvement observed with FoundationGrasp over GraspGPT indicates the necessity of incorporating both semantic and geometric knowledge. The running time of FoundationGrasp during the perception stage is illustrated in Table V. Compared to perception experiments where object and task knowledge in the LaViA-TaskGrasp dataset is pre-generated and pre-processed, it takes an extra 90.40s to prompt foundation models and process the knowledge on the fly in real-robot experiments. FoundationGrasp experiences a significant performance drop from the perception to the action stage. Three typical failure modes are identified: (1) unstable grasp candidates generated by the task-agnostic grasp sampler, (2) motion planning failures, and (3) false rejection of valid candidates. Regarding the third failure mode, the training stage of FoundationGrasp can be conceptualized as constructing an implicit knowledge base in the latent space with a set of objects and tasks. During the inference stage, FoundationGrasp aims to build connections between existing elements in the knowledge base and novel ones outside. However, if novel elements are substantially divergent from the knowledge base, FoundationGrasp tends to make wrong predictions. As illustrated in Figure 8, when instructing the robot to screw with a power drill, the object class power drill shares minimal semantic or geometric similarity to existing elements in the knowledge base, leading to the rejection of all valid candidates. This issue can potentially be resolved by constructing a base element set that is representative enough yet compact.
Phase Perception Stage Total
Segmentation Sampling Prompting Evaluation
Time(s) 7.51 4.32 90.40 1.88 104.11
TABLE V: Running time of FoundationGrasp in real-robot experiments
Multi-Object Task-Oriented Grasping Experiments  To further assess the limits of FoundationGrasp in more complex and realistic environments, we conduct multi-object TOG experiments with three levels of difficulty. The results are presented in Table IV. The success rates remain high (above 70%) in both light and medium clutter settings. This can be attributed to two factors: (1) GroundingDINO+SAM’s ability to accurately detect and segment target objects in less cluttered scenes, and (2) FoundationGrasp’s enhanced geometric understanding, enabling it to interpret the geometric structures of objects and tasks even with partial observations and light occlusions. However, the success rate in heavy clutter drops below 60%. This decrease is expected since heavy occlusions and contacts lead to limited visibility of target objects and complicate the search for collision-free paths to reach them. Nevertheless, the small gap between the grounding and action stages suggests that FoundationGrasp remains effective as long as there is adequate visual observation.